# Private Online Community Detection for Censored Block Models

### v1

@@ 6 Numerical Results
In this section, we present numerical results to assess the performance of our proposed algorithm. The proposed mechanisms are implemented in MATLAB, and the optimization (via SDP method) is done through the CVX solver [15].
6.1 Simulation Study
In the numerical results, we evaluate two performance metrics: 1) the detection delay as defined in (10); and 2) the average run length defined as $\mathbb{E}_{\infty}[\tau]$. In our numerical experiments, we simulate the average detection delay as $\mathbb{E}_{0}[\tau]$ by setting the change-point at time 00.
We set the number of nodes $n=50$, and select the pre- and post-change community label vector $\bm{\sigma}^{\text{pre}}$ and $\bm{\sigma}^{\text{post}}$ such that $\operatorname{Ham}({\bm{\sigma}^{{\text{pre}}}},{\bm{\sigma}^{{\text{post}}}})=2$. The distribution parameter is $p=a\log(n)/n$, and $\zeta=0.1$. The private parameter is set as $\epsilon$. We consider two senarios: (1) Case 1 ($a=5,\epsilon=1.5$) represents the case that condition (19) is satisfied, and we use the smallest window size $w=1$ in such scenario; (2) Case 2 ($a=6,\epsilon=1$) represents the case that condition (19) is violated and the community recovery can be unstable, thus we use multi-window with $w=10$.
We demonstrate the sample trajectories of the detection statistics in Figure 1; we observe that the detection statistics under Case 1 are more stable under the pre-change regime (mostly below zero) and become mostly positive after the change, yielding a relatively small delay as visualized in Figure 2 (Left). We also compare the simulated delay of our proposed method with the lower-bound in Theorem 5.1 in Figure 2 (Left); it can be seen that the delay of our method is close to the lower-bound and matches the lower-bound in the first order. It is worth mentioning that the KL divergence between the post- and pre-change distributions in Case 1 is significant. Thus, the information-theoretic lower bound on the delay is very small, and our proposed method can also obtain a delay of less than four, even for large ARL constraints.
Moreover, we visualize the trajectory for Case 2 in Figure 1 (Right), and the pre-change statistics become more noisy due to the violation of condition (19), while we are still able to detect the change, with a relatively larger delay as shown in Figure 2 (Right), which of course is no longer near-optimal. We want to comment that this shows even if the condition (19) is violated, the proposed detection statistics are still able to detect the change, and we left the more involved theoretical analysis on the choice of window size $w$ and to further improve the detection performance as future work.
@@ 6.1 Simulation Study
Case 1 Case 2
Figure 1: Sample trajectories of the detection statistics.
@@
Case 1 Case 2
Figure 2: Detection Delay v.s Average Run Length.
6.2 Real Case Study
We implemented the online detection algorithm on the worldwide agricultural trade network dataset, which is publicly available at the Food and Agricultural Organization of the United Nations (The dataset is available at https://www.fao.org/faostat/en/#data/TM). We select annual data from year 1991 to 2015. The dataset includes the quantity and value of all food and agricultural products imported/exported annually by all the countries in the world, and it has been previously analyzed in studies such as [34].
We select $n=50$ countries with the most significant importing/exporting relationship to construct the adjacency matrices. For each year, the entry $A_{i,j}$ in the network is set to $A_{i,j}=0$ if there is no import-export relation between two countries, $A_{i,j}=1$ if the total trade column of the top $4$ agricultural products exceeds a pre-specified threshold, and $A_{i,j}=-1$ otherwise. Consequently, each node corresponds to a country, and each edge represents the export-import relationship between countries. We use the data of the year 1991 to estimate the pre-change community label, parameters $p$, and $\zeta$.
Refer to caption Refer to caption Refer to caption Refer to caption
Raw Network $\epsilon=1.5$ $\epsilon=1$ $\epsilon=0.8$
Figure 3: Results on the agricultural trade dataset.
Figure 3 is the plot of the resulting CUSUM detection statistics, which matches well with the change-point 2002 identified by previous studies, such as [34]. As mentioned in [34], the year 2002 was a significant low and turning point in global commodity prices, which likely influenced trade patterns since as countries adjusted their import and export practices in response to changing prices. From the trajectory of detection statistics after applying randomized perturbation with different levels of $\epsilon$ values, we can see the tradeoff between privacy and utility (detection effectiveness): stronger privacy constraints (lower $\epsilon$ values) make it more challenging to detect the change-point.
Furthermore, we also compare the community estimate results using the data in the years 1991 (before the change) and 2003 (after the detected change-point). The estimated community labels for seven countries (United Arab Emirates, Egypt, Hungary, Lebanon, Viet Nam, Saudi Arabia, Romania) changed from $-1$ to $1$. Here, $1$ represents the larger community at both times, indicating that countries such as “Viet Nam” have become more engaged with other leading countries in the global market for agricultural imports and exports after 2002.
7 Conclusion and Discussion
@@
Building on these findings, we identify several avenues for future research. Firstly, we may consider the detection of general changes, including changes in both the community label vector $\bm{\sigma}$ and the parameters $p,\zeta$ of the CBM model. Changes in $p$ and $\zeta$ will represent practical scenarios in which the local connectivity within a community is strengthened or weakened after the change. Secondly, we leave the more involved discussion on the impact of the multi-view stochastic model [39] on our joint detection and estimation scheme, as well as the corresponding impact of the window size $w$, for future investigation. We conjecture that there is an optimal tradeoff between the detection performance and computational/memory efficiency by carefully choosing the window size $w$, especially in the non-asymptotic regime where the network size $n$ is limited. Finally, this work opens numerous opportunities for private community detection problems, and more general differential privacy mechanisms such as the theoretical guarantee under different privacy can be studied for future work.
Appendix A Derivation of the Log-Likelihood and the KL Divergence
@@
For a given graph $G_{t}$ at time $t$ with adjacency matrix ${\mathbf{A}}_{t}$, denote $E_{t,1}:=\#\{i<j:\sigma_{i}A_{t,i,j}\sigma_{j}=1\}$ as the number of edges within the same community with edge label $1$, or in different communities with edge label $-1$; similarly $E_{t,-1}:=\#\{i<j:\sigma_{i}A_{t,i,j}\sigma_{j}=-1\}$. Let $E_{t}=E_{t,1}+E_{t,-1}$ be the total number of edges at time $t$. Then we have the likelihood function is
@@
Therefore the log-likelihood is
$\log p(\mathbf{A};\bm{\sigma},p,\zeta)=E_{t}\log\frac{p}{1-p}+{n\choose 2}\log(1-p)+E_{t,1}\log\frac{1-\zeta}{\zeta}+E_{t}\log\zeta.$
@@
$\frac{1}{2}\bm{\sigma}^{T}\mathbf{A}_{t}\bm{\sigma}=E_{t,1}-E_{t,-1}\Rightarrow 2E_{t,1}-E_{t}=\frac{1}{2}\bm{\sigma}^{T}\mathbf{A}_{t}\bm{\sigma}\Rightarrow E_{t,1}=\frac{1}{4}\bm{\sigma}^{T}\mathbf{A}_{t}\bm{\sigma}+\frac{1}{2}E_{t}.$
Substitute this into the above log-likelihood we obtain
$\log p(\mathbf{A}_{t};\bm{\sigma},p,\zeta)=\left(\frac{1}{4}\log\frac{1-\zeta}{\zeta}\right)\bm{\sigma}^{T}\mathbf{A}_{t}\bm{\sigma}+{n\choose 2}\log(1-p)+E_{t}\log\left(\frac{p}{1-p}\sqrt{\zeta(1-\zeta)}\right).$ (24)
@@
By the log-likelihood function we have that the MLE for $\bm{\sigma}$ is $\hat{\bm{\sigma}}_{\text{MLE}}=\arg\max_{\bm{\sigma}}\{\bm{\sigma}^{T}\mathbf{A}\bm{\sigma}\}$.
KL Divergence (Remark 5.2)
@@
Based on (24), the KL divergence between ${\mathsf{CBM}}(\bm{\sigma},p,\zeta)$ and ${\mathsf{CBM}}(\bm{\sigma}^{\prime},p,\zeta)$ can be calculated as
$I$ $={\mathbb{E}}_{\bm{\sigma}^{\prime}}[\log p({\mathbf{A}};\bm{\sigma}^{\prime},p,\zeta)-\log p({\mathbf{A}};\bm{\sigma},p,\zeta)]$ (25)
$=\left(\frac{1}{2}\log\frac{1-\zeta}{\zeta}\right)p(1-2\zeta){n\choose 2}+{n\choose 2}\log(1-p)+p{n\choose 2}\log\left(\frac{p}{1-p}\sqrt{\zeta(1-\zeta)}\right)$
$\quad\quad-\left(\frac{1}{2}\log\frac{1-\zeta}{\zeta}\right)p(1-2\zeta)\sum_{i<j}\bm{\sigma}_{i}\bm{\sigma}_{j}\bm{\sigma}_{i}^{\prime}\bm{\sigma}_{j}^{\prime}-{n\choose 2}\log(1-p)-p{n\choose 2}\log\left(\frac{p}{1-p}\sqrt{\zeta(1-\zeta)}\right)$
$=\left(\frac{1}{2}\log\frac{1-\zeta}{\zeta}\right)p(1-2\zeta)\left[{n\choose 2}-C_{\bm{\sigma},\bm{\sigma}^{\prime}}\right],$
@@
where $C_{\bm{\sigma},\bm{\sigma}^{\prime}}:=\sum_{i<j}\bm{\sigma}_{i}\bm{\sigma}_{j}\bm{\sigma}_{i}^{\prime}\bm{\sigma}_{j}^{\prime}$ represent the similarity level between the community label vectors $\bm{\sigma}$ and $\bm{\sigma}^{\prime}$ and it satisfies $|C_{\bm{\sigma},\bm{\sigma}^{\prime}}|\leq{n\choose 2}$.
Appendix B Ancillary Lemmas

### v2

@@ V Numerical Results
In this section, we present numerical results to assess the performance of proposed detection frameworks. We emphasize that the examples presented are primarily illustrative. Given the theoretical focus of our work, the simulations are not extensive but aim to demonstrate the fundamental principles and potential applications of our approaches. The proposed method is implemented in MATLAB, and the optimization (via SDP method) is done through the CVX solver [18].
V-A Simulation Study
In this simulation study, we evaluate two performance metrics: 1) the detection delay simulated by assuming the change happens at $\nu=1$; and 2) the average run length $\mathbb{E}_{\infty}[T]$ for the test (stopping time) $T$. We consider a moderate network size of $n=50$, and choose $\bm{\sigma}^{\text{pre}}$ and $\bm{\sigma}^{\text{post}}$ such that $\operatorname{Ham}({\bm{\sigma}^{{\text{pre}}}},{\bm{\sigma}^{{\text{post}}}})=2$, corresponding to a scenario where only a small number of nodes change their community labels. The distribution parameter is set as $p=a\log(n)/n$ for some $a>0$ and $\zeta=0.1$. The private parameter is set as $\epsilon$. We consider two scenarios: (1) Case 1 ($a=5,\epsilon=1.5$) and (2) Case 2 ($a=6,\epsilon=1$). These settings are chosen such that the sufficient condition for exact recovery in Theorem IV.1 is satisfied in Case 1 but violated in Case 2. This allows us to compare detection performances under both a favorable scenario with exact recovery guarantees and a more challenging scenario where exact recovery after perturbation is not assured. It is worth mentioning that the KL divergence between the post- and pre-change distributions after graph perturbation in Case 1 is larger, meaning it is easier to be detected. We also tried multi-window with $w=10$ (see Footnote 2) in the more challenging Case 2 to enhance the detection performance.
We demonstrate the sample trajectories and the trade-off curve of delay vs. ARL, for the detection procedure $T_{\rm L}$ in (8) under LDP settings, in Fig. 5 for Case 1 (left) and Case 2 (right). We observe that the detection statistics under Case 1 are more stable under the pre-change regime (mostly below zero) and become mostly positive after the change, yielding a relatively small delay. On the other hand, the detection statistic for Case 2 is less stable and exhibits several upward trends before the true change point. These could be caused by the stringent privacy constraint, as a smaller $\epsilon$ value in this scenario will lead to significant alterations in data distributions, and the exact recovery sufficient condition (12) is violated.
For completeness, we also include the detection results of the procedure $T_{\rm C}$ in Fig. 6. It is obvious that the performance of the detection procedure here is better than the procedure under LDP settings, as demonstrated by the detection statistics and small detection delay under both cases. This is expected and consistent with our intuition since the LDP constraint is more stringent than the CDP requirement. Under the LDP constraint, we have to privatize the data at the source, which could significantly reduce the divergence between post- and pre-change distributions, especially for small $\epsilon$ values. On the other hand, under the CDP setting, we calculate the detection statistics mostly under the true graph observations. Moreover, the Laplace noise added to the detection statistics is relatively small compared with the recursive statistics $S_{t}$, thus the detection performance could be much better. We highlight that the difference in performance is caused by the distinct nature of LDP and CDP mechanisms, which do not necessarily imply which detection procedure is better than the other.
We further comment on the non-private baseline for change detection, which applies the classic CUSUM test directly to the original data without any privacy protection. As commented in Remark IV.3, the non-private CUSUM method provides the information-theoretic lower bound, $\log\gamma/I_{0}(1+o(1))$, on the detection delay and is guaranteed to incur a smaller delay. In our two scenarios, since the KL divergence between post- and pre-change models $I_{0}$ is relatively large, the corresponding theoretical and empirical delays are both small (approximately one for the ARL values smaller than $10^{4}$). Therefore, we omit this baseline from the plots. Nonetheless, as seen in Fig. 5 and Fig. 6, the detection delays under both LDP and CDP settings (particularly in Case 1) are generally less than 4, indicating highly efficient detection performance even with privacy constraints.
@@ V-A Simulation Study
Figure 7: Comparison of community label recovery errors for the semidefinite relaxation method and the spectral method. Each plot shows the average recovery error over 50 repetitions. The distributional parameters are fixed at $p=0.8$ and $\zeta=0.1$. In the left plot $\epsilon=1.5$ is held constant, while in the right plot, the number of nodes is fixed at $n=50$. The true community label vector $\bm{\sigma}$ consists of half +1+1 entries and half 1-1 entries.
Remark V.1 (Recovery Error).
As a sanity check, we evaluate the community label recovery error of the semidefinite relaxation method and compare it with the spectral method proposed in [12]. Specifically, we apply the privacy mechanism described in Section III-A via graph perturbation, then estimate the underlying community label vector using both the semidefinite relaxation formulation in (3) and the spectral method. The recovery error is measured by the normalized Hamming distance, $\operatorname{Ham}({\bm{\sigma}},{\hat{\bm{\sigma}}})/n$, where $\bm{\sigma}$ is the true community label and $\hat{\bm{\sigma}}$ is the estimated value. As shown in Fig. 7, both methods exhibit comparable recovery performance across various scenarios, with node size $n$ from 20 to 50 and privacy budget $\epsilon$ from 0.6 to 1.5. The parameter values are selected such that the exact recovery condition in (12) is satisfied. Based on this observation, we adopt the semidefinite relaxation method in our work. It is worthwhile mentioning that our proposed detection frameworks are flexible, and they can readily incorporate the spectral method as an alternative for estimating post-change community structures as well.
V-B Real Case Study
We also implement the detection algorithm on a real-world agricultural trade dataset [^3] This dataset is available at https://www.fao.org/faostat/en/#data/TM. and a U.S. air transportation network dataset [^4] This dataset is available at here..
Agricultural Trade Dataset
The dataset includes the quantity and value of all food and agricultural products imported/exported annually by all the countries in the world. We select annual data from the years 1991 to 2015 and $n=50$ countries with the most significant importing/exporting relationship. For each country pair $(i,j)$, we let $A_{i,j}=0$ if there is no import-export relation between them, $A_{i,j}=1$ if the total trade column of the top $4$ agricultural products exceeds a pre-specified threshold, and $A_{i,j}=-1$ otherwise. Consequently, each node corresponds to a country, and each edge represents the export-import relationship between countries. We use the data of the year 1991 to estimate the pre-change community label, parameters $p$, and $\zeta$.
Refer to caption Refer to caption Refer to caption Refer to caption
Raw Network ($\epsilon=+\infty$) $\epsilon=1.5$ $\epsilon=1$ $\epsilon=0.8$
Figure 8: Trajectory of the detection statistics under LDP on the agricultural trade dataset, for the raw network (no perturbation), the privatized/perturbed network with $\epsilon=1.5$, 1, and 0.8, representing different privacy levels.
Since private detection under LDP settings is more challenging than in CDP for the same privacy parameter $\epsilon$, we visualize the detection statistics from (7) under LDP in Fig. 8. The detection statistics successfully identify the change-point in 2002, as noted in previous studies [38]. This year 2002 marked a significant turning point in global commodity prices, which likely influenced trade patterns of various countries [38]. Furthermore, from the trajectory of detection statistics after applying randomized perturbation with different levels of $\epsilon$ values, we can see the tradeoff between privacy and utility (detection effectiveness): stronger privacy constraints (lower $\epsilon$ values) make it more challenging to detect the change.
U.S. Air Transportation Network Dataset
We use monthly data from February 2018 to July 2022 (54 months). Each node represents an airport, and each edge encodes the presence of direct flights between two airports. We select $n=50$ airports with the highest combined numbers of departing and arriving flights. For each airport pair $(i,j)$, we let $A_{i,j}=0$ if there is no direct flight recorded, $A_{i,j}=1$ if the number of airlines operating direct flights between them exceeds a pre-specified threshold, and $A_{i,j}=-1$ otherwise. Consequently, the resulting graph represents the density of airline connectivity among major U.S. airports. Data from February 2016 to June 2018 are used as the historical data for estimating the pre-change community structure and parameters, with the rest being the dataset for conducting change-point detection.
In this real data example, it is important to note that the distribution parameters $p$ and $\zeta$ may also change over time. Therefore, in addition to computing detection statistics under the assumption of constant $p$ and $\zeta$, we also implement a version that allows these parameters to change. Specifically, we estimate their post-change values jointly with $\bm{\hat{\sigma}}_{t-1}$ using the maximum likelihood approach detailed in Appendix A-B.
@@ V-B Real Case Study
Raw Network ($\epsilon=+\infty$); Assume unchanged $p,\zeta$ Raw Network ($\epsilon=+\infty$); Assume changed $p,\zeta$
Refer to caption Refer to caption
$\epsilon=1.5$; Assume unchanged $p,\zeta$ $\epsilon=0.8$; Assume changed $p,\zeta$
Figure 9: Trajectory of the detection statistics under LDP on the U.S. air transportation dataset, for the raw network (no perturbation, corresponding to $\epsilon=\infty$) and the privatized/perturbed network (with $\epsilon=1.5$). For each case, we report two versions of the detection statistic: one assuming the parameters $p$ and $\zeta$ remain unchanged after the change-point, and another allowing them to change, in which case the post-change values of $p$ and $\zeta$ are also estimated during detection.
We evaluate the more challenging private change detection under local differential privacy (LDP) constraints. Fig. 9 shows the trajectory of the detection statistics. The statistic successfully identifies a change-point in March 2020, coinciding with the disruption in air travel caused by the COVID-19 outbreak, also consistent with findings in previous studies [38]. From the trajectory of detection statistics after applying randomized perturbation, we can see that the detection becomes less effective due to privacy requirements, as evidenced in the slower increase in detection statistics. Furthermore, the detection statistics that take into account the potential change in $p$ and $\zeta$ improve detection performance in both raw and privatized settings, as reflected by the faster growth of the corresponding detection statistics.
VI Conclusion and Discussion
@@
This work has studied online change detection of dynamic community structures under the censored block model. Detection algorithms have been presented under two differential privacy settings. The proposed methods are applicable to a wide range of network types and sizes. Beyond network data, the detection framework can also be extended to other data types, such as spatio-temporal data, where spatial dependencies are captured by an underlying connectivity graph. Building on these findings, we can identify several avenues for future research. Firstly, we may consider the detection of general changes, including changes in both the community label vector $\bm{\sigma}$ and the parameters $p,\zeta$ of the CBM model. Changes in $p$ and $\zeta$ will represent practical scenarios in which the local connectivity within a community is strengthened or weakened after the change. Secondly, we leave the more involved discussion on the impact of the multi-view stochastic model [42] on our joint detection and estimation scheme, as well as the corresponding impact of the window size $w$, for future investigation. We conjecture that there is an optimal tradeoff between the detection performance and computational/memory efficiency by carefully choosing the window size, especially in the non-asymptotic regime where the network size $n$ is limited. Finally, this work opens numerous opportunities for exploring more advanced differential privacy mechanisms to further enhance detection performance.
Appendix A More Technical Details and Discussions
A-A Tradeoff between Privacy and Separation Condition for Graph Perturbation
From Theorem IV.1, we have a sufficient condition for exact recovery being
$a\big(\sqrt{1-\zeta}-\sqrt{\zeta}\big)^{2}>\frac{1}{\alpha}\times\left(\frac{e^{\epsilon}+1}{e^{\epsilon}-1}\right),$
where $\alpha=1-\frac{1}{\sqrt{n}}$. As can be shown in Fig. 10, the analysis of the lower bounds on the parameter $a$ for $\alpha$, specifically $\alpha=1-\frac{1}{\sqrt{n}}=1-o(1)$. This functional form is desirable as it allows for more flexibility and potentially less stringent requirements on the parameter $a$ while maintaining the desired privacy level defined by $\epsilon=c\log(n)$. Thus, in scenarios where achieving an optimal balance between privacy and parameter constraints is crucial, opting for $\alpha=1-o(1)$ would be a more strategic choice.
A-B Derivation of the Log-Likelihood and the KL Divergence
@@
For a given graph $G_{t}$ at time $t$ with adjacency matrix ${\mathbf{A}}_{t}$, denote $E_{t,1}:=\#\{i<j:\sigma_{i}A_{t,i,j}\sigma_{j}=1\}$ as the number of edges within the same community with edge label $1$, or in different communities with edge label 1-1; similarly $E_{t,-1}:=\#\{i<j:\sigma_{i}A_{t,i,j}\sigma_{j}=-1\}$. Let $E_{t}=E_{t,1}+E_{t,-1}$ be the total number of edges at time $t$. Then we have the log-likelihood function is
@@
$\frac{1}{2}\bm{\sigma}^{T}\mathbf{A}_{t}\bm{\sigma}=E_{t,1}-E_{t,-1}=2E_{t,1}-E_{t}$
$\Rightarrow$ $E_{t,1}=\frac{1}{4}\bm{\sigma}^{T}\mathbf{A}_{t}\bm{\sigma}+\frac{1}{2}E_{t}.$
Substitute this into the log-likelihood we obtain
$\log p(\mathbf{A}_{t};\bm{\sigma},p,\zeta)=\left(\frac{1}{4}\log\frac{1-\zeta}{\zeta}\right)\bm{\sigma}^{T}\mathbf{A}_{t}\bm{\sigma}$
$+{n\choose 2}\log(1-p)+E_{t}\log\left(\frac{p}{1-p}\sqrt{\zeta(1-\zeta)}\right).$ (22)
Maximum Likelihood Estimator (MLE)
@@
By the log-likelihood function (A-B), we have that the MLE for $\bm{\sigma}$ is $\hat{\bm{\sigma}}_{\text{MLE}}=\arg\max_{\bm{\sigma}}\{\bm{\sigma}^{T}\mathbf{A}\bm{\sigma}\}$. And the MLE for $p$ and $\zeta$ are $\hat{p}_{\text{MLE}}=E_{t}/{n\choose 2}$ and $\hat{\zeta}_{\text{MLE}}=\frac{1}{2}-\frac{\hat{\bm{\sigma}}_{\text{MLE}}^{T}\mathbf{A}\hat{\bm{\sigma}}_{\text{MLE}}}{4E_{t}}$, respectively.
KL Divergence
@@
Based on (A-B), the KL divergence between ${\mathsf{CBM}}(\bm{\sigma},p,\zeta)$ and ${\mathsf{CBM}}(\bm{\sigma}^{\prime},p,\zeta)$ can be calculated as
$I$ $={\mathbb{E}}_{\bm{\sigma}^{\prime},p,\zeta}[\log p({\mathbf{A}};\bm{\sigma}^{\prime},p,\zeta)-\log p({\mathbf{A}};\bm{\sigma},p,\zeta)]$
$=\left(\frac{1}{2}\log\frac{1-\zeta}{\zeta}\right)p(1-2\zeta)\left[{n\choose 2}-C_{\bm{\sigma},\bm{\sigma}^{\prime}}\right],$ (23)
@@
where $C_{\bm{\sigma},\bm{\sigma}^{\prime}}:=\sum_{i<j}{\sigma}_{i}{\sigma}_{j}{\sigma}_{i}^{\prime}{\sigma}_{j}^{\prime}$ represent the similarity level between the community label vectors $\bm{\sigma}$ and $\bm{\sigma}^{\prime}$ and it satisfies $|C_{\bm{\sigma},\bm{\sigma}^{\prime}}|\leq{n\choose 2}$.
A-C Analysis of the Graph Perturbation Mechanism
In this subsection, we provide analysis for the Ternary Randomized Response Mechanism used in Section III-A. In this mechanism, when $A_{i,j}=x$ where $x\in\{0,+1,-1\}$, the mechanism outputs $\tilde{A}_{i,j}=y$, where $y\in\{0,+1,-1\}$ with probability $q_{x}(y)\triangleq\operatorname{Pr}(\tilde{{\mathbf{A}}}_{i,j}=y|{\mathbf{A}}_{i,j}=x)$. We set $q_{x}(y)=c_{2}$, $\forall x\neq y$, and $q_{x}(y)=c_{1}=1-2c_{2}$, $\forall x=y$. We denote $\tilde{p}$ as the distribution of each edge $\tilde{A}_{i,j}$ for the pair $(i,j)$ such that $\bm{\sigma}_{i}=\bm{\sigma}_{j}$ (in the same community), and denote $\tilde{q}$ as the distribution of each edge $\tilde{A}_{i,j}$ for the pair $(i,j)$ such that $\bm{\sigma}_{i}\neq\bm{\sigma}_{j}$ (in different communities). We have the following:
$\tilde{p}({+1})$ $=c_{1}p(1-\zeta)+c_{2}p\zeta+c_{2}(1-p),$
$\tilde{p}({-1})$ $=c_{2}p(1-\zeta)+c_{1}p\zeta+c_{2}(1-p),$
$\tilde{p}({0})$ $=c_{2}p(1-\zeta)+c_{2}p\zeta+c_{1}(1-p),$
$\tilde{q}({+1})$ $=c_{1}p\zeta+c_{2}p(1-\zeta)+c_{2}(1-p),$
$\tilde{q}({-1})$ $=c_{2}p\zeta+c_{1}p(1-\zeta)+c_{2}(1-p),$
$\tilde{q}({0})$ $=c_{2}p(1-\zeta)+c_{2}p\zeta+c_{1}(1-p).$
[…]
$\tilde{p}$ $=1-c_{1}(1-p)-c_{2}p=2c_{2}+p(1-3c_{2}),$
$\tilde{\zeta}$ $=\frac{c_{2}(1-p\zeta)+c_{1}p\zeta}{1-c_{1}(1-p)-c_{2}p}=\frac{c_{2}+p\zeta(1-3c_{2})}{2c_{2}+p(1-3c_{2})}.$ (24)
We further analyze the change in the KL divergence between the pre- and post-change distributions after the perturbation. The KL divergence for the raw data distribution (before perturbation) can be derived as (see (A-B)),
$I(\bm{\sigma}^{\text{pre}};\bm{\sigma}^{\text{post}})=\frac{1}{2}\log\frac{1-\zeta}{\zeta}\cdot p(1-2\zeta)\left[{n\choose 2}-C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}}\right].$ (25)
After the perturbation, we have similarly
$\tilde{I}(\bm{\sigma}^{\text{pre}};\bm{\sigma}^{\text{post}})=\frac{1}{2}\log\frac{1-\tilde{\zeta}}{\tilde{\zeta}}\cdot\tilde{p}(1-2\tilde{\zeta})\left[{n\choose 2}-C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}}\right].$ (26)
Substitute the parameter $\tilde{p}$ and $\tilde{\zeta}$ in (A-C), and using the setting that $c_{1}=1-2c_{2}$, we have
[…]
$=$ $\left(\log\frac{c_{2}+p(1-\zeta)(1-3c_{2})}{c_{2}+p\zeta(1-3c_{2})}\right)\cdot p(1-2\zeta)(1-3c_{2}).$ (27)
We note that $\log\frac{c_{2}+p(1-\zeta)(1-3c_{2})}{c_{2}+p\zeta(1-3c_{2})}$ is a positive and decreasing function of $c_{2}$ when $\zeta<\frac{1}{2}$ and a negative and increasing function when $\zeta>\frac{1}{2}$, the same is the function $p(1-2\zeta)(1-3c_{2})$. Moreover, when $c_{2}=0$ we have the KL divergence remaining unchanged. Combining these together, we have (A-C) is a nonnegative and decreasing function of $c_{2}$, with the largest value obtained when $c_{2}=0$ and the largest value equal to $I(\bm{\sigma}^{\text{pre}};\bm{\sigma}^{\text{post}})$. This demonstrates the tradeoff between detection and privacy. This being said, for a given privacy constraint, we may select the smallest value of $c_{2}$ value that satisfies this constraint; such choice ensures the best detection performance after the perturbation.
Note that, under $\epsilon$-edge LDP, we require $c_{2}/c_{1}\in[e^{-\epsilon},e^{\epsilon}]$, which yields the smallest possible value for $c_{2}$ being $c_{2}=\frac{1}{e^{\epsilon}+2}$, which is the value we choose in III-A.
Appendix B Fundamentals of Technical Analysis
B-A Ancillary Lemmas